4. A spherical balloon of 0.5 m diameter contains air at a pressure of 500 kPa. The diameter increases to 0.55m in a reversible process during which pressure is proportional to diameter. Determine the work done by the air in the balloon during this process. Also calculate the final pressure.


1
Expert's answer
2021-09-06T00:49:55-0400

From the condition of the process P = CD

Where p is the pressure (kPa)

C is a constant

D is the diameter (m) From the initial condition

500 = C * 0.5

C = 500 / 0.5 = 1000 Therefore

P = 1000 D

1W2 ="\int" 1000DdV


To substitute dV in terms of dD Consider the volume of the sphere

"V=\pi /6 (D)" 3

Now,

"=500*\pi/4" (0.554-0.504)

"=11.39 KJ"


Final pressure


P2 = 100

D2 = 1000 * 0.55

= 550 kPa